# Deep Instruction Tuning for Segment Anything Model

### v1

@@ 3.2 End-to-end Instruction Tuning
One straightforward solution for SAM is to directly feed the prompt words to its ViT-based image encoder. Due to the flexibility of Transformer architecture, the text features can be also regarded as additional input tokens for feature transformations. This attempt is also validated in previous VL works like ViLT (Kim et al., 2021), and has a similar principle with the popular prompt tuning (Jia et al., 2022).
@@
Here, we term this DIT solution as the end-to-end tuning, abbreviated as E-DIT. Concretely, given a text description $T$, we first use the language encoder, i.e., BERT (Devlin et al., 2018), to extract their representations, denoted as $\mathbf{F}_{t}\in\mathbb{R}^{l\times d}$, where $l$ denotes the length of description. The visual features after the patch embedding layer is denoted as $\textbf{F}_{v}^{0}\in\mathbb{R}^{m\times d}$. Here, $m=HW/P^{2}$ is the number of patches, where $H\times W$ is the image resolution and $P^{2}$ is the patch size. Afterwards, we use a linear layer to project $\mathbf{F}_{t}$ onto the visual semantic space of SAM, and obtain $\mathbf{F}_{t}^{0}$:
$\mathbf{F}_{t}^{0}=\mathbf{F}_{t}\textbf{W}_{0}+b_{0},$ (4)
@@
These text tokens and image patches are then processed by the following Transformer layers of the encoder for both image feature learning and cross-modal fusions. The output image features $\mathbf{F}_{v}^{n}$ alone with the text features $\mathbf{F}_{t}$ are fed to the lightweight mask decoder. In this case, E-DIT is formulated by
$\mathbf{F}_{m}^{0}=[\mathbf{F}_{t}^{0};\mathbf{F}_{v}^{0}],$ (5)
$\mathbf{F}_{m}^{i+1}=L_{i}(\mathbf{F}_{m}^{i}),$ (6)
$M^{\prime}=\mathcal{F}_{d}(\mathbf{F}_{m}^{n},\mathbf{F}_{t}\mathbf{W}_{o}).$ (7)
@@
Considering the case of freezing SAM’s backbone, Eq.4 only adds a few parameters for semantic projection. Meanwhile, when the hidden Transformer layers are frozen, the inserted text tokens do not fit well into visual space. For the output of each layer, we also add a linear projection to facilitate text adaption, as shown in Fig. 3.
Table 1: Comparison of different text-instructed schemes on RefCOCO. Default refers to the default lightweight fusion scheme of SAM. “Add dec layers” is to increasing the depth of the mask decoder. * denotes that the image backbone is frozen, while the text and mask decoders are tuned. The best and second best performances are in bold and underlined.
@@
Table 2: The combination with conventional prompt tuning. Here, dynamic refers to the projected text words changed according to different examples, i.e., the setting of DITs, and static denotes the learnbale tokens.
@@
3.3 Layer-wise Instruction Tuning
According to Wu et al. (2023b), it is often difficult for additionally inserted tokens to fully interact with the existing sequences in the well learned semantic space, especially under the end-to-end transformations. We consider the layer-wise text instruction tuning for SAM, demoted as L-DIT.
Concretely, we insert the text prompts to the sub-semantic space of each Transformer layer of SAM’ encoder, thereby mitigating the modality gap for better tuning. Given the text features $\mathbf{F}_{t}$, we linearly project them to each layer of the image encoder, i.e., $\mathbf{F}_{t}^{i}$, and the input sequence is defined by
$\mathbf{F}_{m}^{i}=[\mathbf{F}_{t}^{i};\mathbf{F}_{v}^{i}],$ (8)
@@ 3.3 Layer-wise Instruction Tuning
Similar to Eq. 5-7, the final output visual tokens $\mathbf{F}_{v}^{k}$ as well as the projected text features $\mathbf{F}_{t}^{\prime}$ are fed to the lightweight decoder for mask prediction. Compared with E-DIT, this layer-wise scheme can better adapt text prompts to the subspaces of SAM’s encoder, thereby improving the efficiency of text instruction tuning.
Table 3: The impact of freezing the visual and text encoders. Here, ✓means that the encoder is frozen, while × is updated.
@@
× End-to-end 76.51 66.00 23.19 62.96 66.87 78.89 68.41 21.31 63.74 67.83 72.90 60.01 25.39 60.43 64.63
× Layer-wise 81.22 70.21 24.31 67.50 69.97 83.89 74.66 24.36 69.55 72.44 77.26 65.44 28.26 64.77 68.12
× × End-to-end 81.78 73.07 28.27 68.07 71.46 85.28 77.17 27.56 70.81 73.90 78.02 67.60 30.20 65.58 69.36
× × Layer-wise 85.68 76.61 29.89 71.98 74.73 88.06 79.17 25.94 74.51 75.62 81.28 69.11 30.95 68.77 71.21
Table 4: The impact of different text words injections for L-DIT. CrossAtt and FFN refer to mutli-head cross attention and feed-forward network (Vaswani et al., 2017), respectively. Linear projection is the setting of L-DIT.
@@
Table 5: Comparison with the state-of-the-art methods on three RIS datasets. Here, Scratch refers to the models only tuned with RIS examples of these datasets, while Pre-trained denotes the use of a large number of visual grounding examples (Krishna et al., 2017) for pre-training. The metric used is oIoU.
Method Visual Backbone Textual Encoder RefCOCO RefCOCO+ RefCOCOg
@@
MCN (Luo et al., 2020b) Darknet53 bi-GRU 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
EFN (Feng et al., 2021) ResNet101 bi-GRU 62.76 65.69 59.67 51.50 55.24 43.01 - -
BUSNet (Yang et al., 2021) DeepLab-R101 Self-Att 63.27 66.41 61.39 51.76 56.87 44.13 - -
CGAN (Luo et al., 2020a) DeepLab-R101 bi-GRU 64.86 68.04 62.07 51.03 55.51 44.06 51.01 51.69
ISFP (Liu et al., 2022) Darknet53 bi-GRU 65.19 68.49 62.73 52.70 56.77 46.39 52.67 53.00
LST (Jing et al., 2021) Darknet53 bi-GRU 65.43 67.76 63.08 54.21 58.32 48.02 54.40 54.25
ReSTR (Kim et al., 2022) ViT-B Transformer 67.22 69.30 64.45 55.78 60.44 48.27 - -
LAVT (Yang et al., 2022) Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
VLT (Ding et al., 2022) Swin-B BERT 72.96 75.96 69.60 63.53 68.43 56.92 63.49 66.22
VDP (Zhao et al., 2023) VQGAN CLIP 73.25 - - 62.69 - - 61.96 -
ReLA (Liu et al., 2023b) Swin-B BERT 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
SADLR (Yang et al., 2023) Swin-B BERT 74.24 76.25 70.06 64.28 69.09 55.19 63.60 63.56
@@
MaLT (Li et al., 2021) ViLT BERT 70.13 71.71 66.92 62.23 65.92 56.06 62.45 62.87
CRIS (Wang et al., 2022) CLIP-R101 CLIP 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
ETRIS (Xu et al., 2023b) CLIP-R101 CLIP 71.06 74.11 66.66 62.23 68.51 52.79 60.28 60.42
SeqTR (Zhu et al., 2022) Darknet53 bi-GRU 71.70 73.31 69.82 63.04 66.73 58.97 64.69 65.74
SEEM (Zou et al., 2024) DaViT-d5 Florence - - - - - - 65.70 -
LISA-7B (Lai et al., 2023) ViT-H LLaVA-7B 74.90 79.10 72.30 65.10 70.80 58.10 67.90 70.60
PolyFormer (Liu et al., 2023d) Swin-L BERT 75.96 78.29 73.25 69.33 74.56 61.87 69.20 70.19
@@
4 Experiments
To validate the proposed DITs, we conduct extensive experiments on three benchmark datasets of RIS, i.e., RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016), and also compare L-DIT with a bunch of state-of-the-art (SOTA) methods in RIS (Zhu et al., 2022; Yang et al., 2022; Liu et al., 2023b, d).
4.1 Datasets
4.2 metrics
We use the overall intersection-over-union (oIoU), the mean intersection-over-union (mIoU), and precision at threshold of 0.5, 0.7, and 0.9 as evaluation metrics. The oIoU is measured by the ratio of the total area of intersection to the total area of union across all test samples. The metric favors large objects over small ones. The mIoU is calculated as the average of the IoU values between predictions and the ground truth for all test samples. This measure is applied uniformly to both large and small objects.
4.3 Implementation Details
For visual backbone, we use ViT-B/16, initialized from pre-trained weights of SAM. BERT with 12 layers and a hidden dimension of 768 is used as language encoder. We keep the aspect ratio and resize the image to 512 x 512. Following this, we apply color augmentation, Gaussian blur and random horizontal flipping to augment the data. To binarize the prediction of RIS, we set a threshold of 0.35. The optimizer is Adam with a learning rate of 1e-4, which is multiplied by a decay factor of 0.2 at the 30th and the 35th, and the batch size is set to 64. The warm-up epoch is configured to 3. The training process spans 40 epochs.
4.4 Quantitative analysis
4.4.1 Comparison with the default scheme
@@
In Tab. LABEL:Ablation_of_Tuning_method, we first compare our deep instruction tuning (DIT) methods with the default fusion scheme of SAM and its alternatives, i.e., adding more layers to the mask decoder. We can first see that the default fusion scheme performs much worse than DITs on RefCOCO under the setting of either fully tuning or freezing backbone, suggesting that the default lightweight decoder is insufficient in image-text interactions. With more decoding layers, this issue can be alleviated to some extent, but the performance still lags behind DITs especially when the image encoder frozen. To explain, the additional fusion layers can enhance the cross-modal interactions, but also undermines the powerful segmentation capability of SAM pre-trained on large amounts of data. In contrast, without modifying the main structure of SAM, our DIT methods can improve the performance greatly. Another interesting finding from Tab. LABEL:Ablation_of_Tuning_method is that the advantage of layer-wise DIT is more obvious under the setting of freezing backbone. This result confirms our assumption that the layer-wise projection can better help to project text tokens onto the semantic space of SAM. Overall, Tab. LABEL:Ablation_of_Tuning_method well validates the effectiveness of DITs for text-instructed SAM.
4.4.2 Ablation Study
@@
Combination with prompt tuning. Tab. LABEL:framwork present the results of combing DITs with conventional prompt tuning (Zhou et al., 2022a, b; Wu et al., 2023b), i.e., adding learnable tokens to the text prompts. Under end-to-end tuning, the addition of learnable tokens can improve performance to a certain extent, mainly by increasing the parameter capacity of DIT for better text tuning. However, these prompt tokens are not helpful for layer-wise DIT, on the contrary, reducing performance greatly. This result may suggest that the text words are well projected onto each layer with independent projectors. The addition of parameterized tokens somewhat declines their semantics.

### v2

@@ 3.2. End-to-end Deep Instruction Tuning
One straightforward solution for SAM is to directly feed the instruction words to its ViT-based image encoder. Due to the flexibility of Transformer architecture, the text features can be also regarded as additional input tokens for feature transformations. This attempt is also validated in previous VL works like ViLT (Kim et al., 2021), and has a similar process with the popular prompt tuning (Jia et al., 2022).
@@
Here, we term this DIT solution as the end-to-end tuning, abbreviated as E-DIT. Concretely, given a text instruction $T$, we first use the language encoder, i.e., BERT (Devlin et al., 2018), to extract its representations, denoted as $\mathbf{F}_{t}\in\mathbb{R}^{l\times d}$, where $l$ denotes the length of instruction. The visual features after the patch embedding layer are denoted as $\textbf{F}_{v}^{0}\in\mathbb{R}^{m\times d}$. Here, $m=H\times W/P^{2}$ is the number of patches, where $H\times W$ is the image resolution and $P^{2}$ is the patch size. Afterwards, we use a linear layer to project $\mathbf{F}_{t}$ onto the visual semantic space of SAM, and obtain $\mathbf{F}_{t}^{0}$ via
(4) $\mathbf{F}_{t}^{0}=\mathbf{F}_{t}\textbf{W}_{0}+b_{0},$
@@
These text tokens and image patches are then processed by the following Transformer layers of the encoder for both image feature learning and cross-modal fusion. The output image features $\mathbf{F}_{v}^{n}$ alone with the text features $\mathbf{F}_{t}$ are fed to the lightweight mask decoder. In this case, E-DIT is formulated by
(5) $\mathbf{F}_{m}^{0}=[\mathbf{F}_{t}^{0};\mathbf{F}_{v}^{0}],$
(6) $\mathbf{F}_{m}^{i+1}=L_{i}(\mathbf{F}_{m}^{i}),$
(7) $M^{\prime}=\mathcal{F}_{d}(\mathbf{F}_{v}^{n},\mathbf{F}_{t}\mathbf{W}_{o}).$
@@
Considering the case of freezing SAM’s backbone, Eq. 4 only adds a few parameters for semantic projection. When the hidden Transformer layers are frozen, the inserted text tokens do not fit well into visual space. For the output of each layer, we also add a linear projection to facilitate text adaption.
Table 1. Comparison of different text-instructed schemes on RefCOCO. Default refers to the default lightweight fusion scheme of SAM. Add dec layers is to increasing the depth of the mask decoder. * denotes that the image backbone is frozen, while the text and mask decoders are tuned. The best and second best performances are in bold and underlined.
@@
Table 2. The combination with conventional prompt tuning. Here, dynamic refers to the projected text words changed according to different examples, i.e., the setting of DITs, and static denotes the learnbale tokens.
@@
3.3. Layer-wise Deep Instruction Tuning
According to Wu et al. (Wu et al., 2023b), it is often difficult for additionally inserted tokens to fully interact with the existing feature sequences in the well learned semantic space, especially under the end-to-end manner. In this case, we also propose an effective layer-wise instruction tuning method for SAM, demoted as L-DIT.
Concretely, we project the text tokens onto the sub-semantic space of each Transformer layer of SAM’ encoder, thereby mitigating the modality gap for better tuning. Given the text features $\mathbf{F}_{t}$, we linearly insert them into each layer of the image encoder, i.e., $\mathbf{F}_{t}^{i}$, and the input sequence is defined by
(8) $\mathbf{F}_{m}^{i}=[\mathbf{F}_{t}^{i};\mathbf{F}_{v}^{i}],$
@@ 3.3. Layer-wise Deep Instruction Tuning
Similar to Eq. 5-7, the final output visual tokens $\mathbf{F}_{v}^{k}$ as well as the projected text features $\mathbf{F}_{t}^{\prime}$ are fed to the lightweight decoder for mask prediction. Compared with E-DIT, this layer-wise scheme can better project text tokens onto the subspaces of SAM’s encoder, thereby improving the efficiency of text instruction tuning.
4. Experiments
To validate the proposed DITs, we conduct extensive experiments on three highly competitive benchmarks of RIS, i.e., RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016), and also compare L-DIT with a bunch of state-of-the-art (SOTA) methods (Zhu et al., 2022; Yang et al., 2022; Liu et al., 2023b, c) in RIS.
4.1. Datasets and Metrics
Metrics. We use the overall intersection-over-union (oIoU), the mean intersection-over-union (mIoU), and precision at threshold of 0.5, 0.7, and 0.9 as evaluation metrics. The oIoU is measured by the ratio of the total area of intersection to the total area of union across all test samples. The metric favors large objects over small ones. The mIoU is calculated as the average of the IoU values between predictions and the ground truth for all test samples. This measure is applied uniformly to both large and small objects.
Table 3. The impact of freezing the visual and text encoders. Here, ✓means that the encoder is frozen, while ✗is updated.
@@ 4.1. Datasets and Metrics
End-to-end 76.51 66.00 23.19 62.96 66.87 78.89 68.41 21.31 63.74 67.83 72.90 60.01 25.39 60.43 64.63
Layer-wise 81.22 70.21 24.31 67.50 69.97 83.89 74.66 24.36 69.55 72.44 77.26 65.44 28.26 64.77 68.12
End-to-end 81.78 73.07 28.27 68.07 71.46 85.28 77.17 27.56 70.81 73.90 78.02 67.60 30.20 65.58 69.36
Layer-wise 85.68 76.61 29.89 71.98 74.73 88.06 79.17 25.94 74.51 75.62 81.28 69.11 30.95 68.77 71.21
Table 4. The impact of different text word injections for L-DIT. CrossAtt and FFN refer to mutli-head cross attention and feed-forward network (Vaswani et al., 2017), respectively. Linear projection is the setting of L-DIT.
@@
Table 5. The impact of different text encoders for L-DIT.
Text encoder val testA testB
oIoU mIoU oIoU mIoU oIoU mIoU
BERT 67.50 69.97 69.55 72.44 64.77 68.12
CLIP 67.19 69.82 68.99 71.60 64.88 67.67
Table 6. The impact of tuning on SAM. Here, means not tuning, and Mix denotes the joint training with point, box and text [^4] Boxes and points are only processed in the default decoder.. The metric used is oIoU.
Setting Train Prompt val testA testB Inference
Default SAM Point 75.84 78.11 75.38 119ms
Default SAM Box 75.32 75.38 76.60 120ms
Default SAM Text Text 60.74 62.95 56.69 123ms
Layer-wise DIT Mix Point 85.16 85.59 84.51 123ms
Layer-wise DIT Mix Box 87.47 87.43 87.74 123ms
Layer-wise DIT Mix Text 69.35 71.04 66.76 123ms
Layer-wise DIT Text Text 71.98 74.51 68.77 123ms
4.2. Experimental Setups
For visual backbone, we use ViT-B/16 initialized from pre-trained weights of SAM. BERT with 12 layers and a hidden dimension of 768 is used as language encoder. We keep the aspect ratio and resize the image to 512 $\times$ 512. Following previous works (Luo et al., 2020a; Zhu et al., 2022), we apply color augmentation, Gaussian blur and random horizontal flipping to augment the data. To binarize the prediction of RIS, we set a threshold of 0.35. The optimizer is Adam (Kingma and Ba, 2014) with a learning rate of 1e-4, which is multiplied by a decay factor of 0.2 at the 30-th and the 35-th, and the batch size is set to 64. The warm-up epoch is configured to 3. The training process spans 40 epochs. More details can refer to our anonymously released project.
Table 7. Comparison with the state-of-the-art methods on three RIS datasets. Here, Scratch refers to only using the RIS examples of these datasets for training, while Pre-trained denotes the use of a large number of visual grounding examples (Krishna et al., 2017) for pre-training. The metric used is oIoU.
Method Visual Backbone Textual Encoder Pretraining Data RefCOCO RefCOCO+ RefCOCOg
@@ 4.2. Experimental Setups
MCN(CVPR20) (Luo et al., 2020b) Darknet53 bi-GRU 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
EFN(CVPR21) (Feng et al., 2021) ResNet101 bi-GRU 62.76 65.69 59.67 51.50 55.24 43.01 - -
BUSNet(CVPR21) (Yang et al., 2021) DeepLab-R101 Self-Att 63.27 66.41 61.39 51.76 56.87 44.13 - -
CGAN(MM20) (Luo et al., 2020a) DeepLab-R101 bi-GRU 64.86 68.04 62.07 51.03 55.51 44.06 51.01 51.69
ISFP(MM22) (Liu et al., 2022) Darknet53 bi-GRU 65.19 68.49 62.73 52.70 56.77 46.39 52.67 53.00
LST(CVPR21) (Jing et al., 2021) Darknet53 bi-GRU 65.43 67.76 63.08 54.21 58.32 48.02 54.40 54.25
ReSTR(CVPR22)(Kim et al., 2022) ViT-B Transformer 67.22 69.30 64.45 55.78 60.44 48.27 - -
LAVT(CVPR22) (Yang et al., 2022) Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
VLT(TPAMI22) (Ding et al., 2022) Swin-B BERT 72.96 75.96 69.60 63.53 68.43 56.92 63.49 66.22
VDP(ICCV23) (Zhao et al., 2023) VQGAN CLIP 73.25 - - 62.69 - - 61.96 -
ReLA(CVPR23) (Liu et al., 2023b) Swin-B BERT 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
SADLR(AAAI23) (Yang et al., 2023) Swin-B BERT 74.24 76.25 70.06 64.28 69.09 55.19 63.60 63.56
CARIS(MM23) (Liu et al., 2023f) Swin-B BERT 75.14 78.10 71.75 66.63 72.04 58.51 64.31 65.82
@@
MaLT(arXiv21) (Li et al., 2021) ViLT BERT 3.88M 70.13 71.71 66.92 62.23 65.92 56.06 62.45 62.87
CRIS(CVPR22) (Wang et al., 2022) CLIP-R101 CLIP 3.88M 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
ETRIS(ICCV23) (Xu et al., 2023a) CLIP-R101 CLIP 3.88M 71.06 74.11 66.66 62.23 68.51 52.79 60.28 60.42
SeqTR(ECCV22) (Zhu et al., 2022) Darknet53 bi-GRU 6.10M 71.70 73.31 69.82 63.04 66.73 58.97 64.69 65.74
SEEM(NeurIPS23) (Zou et al., 2023) DaViT-d5 Florence 5.00M - - - - - - 65.70 -
LISA-7B(arXiv23) (Lai et al., 2023) ViT-H LLaVA-7B 18.16M 74.90 79.10 72.30 65.10 70.80 58.10 67.90 70.60
PolyFormer(CVPR23) (Liu et al., 2023c) Swin-L BERT 5.70M 75.96 78.29 73.25 69.33 74.56 61.87 69.20 70.19
@@
4.3. Quantitative Analysis
4.3.1. Comparison with the default scheme
@@
In Tab. LABEL:Ablation_of_Tuning_method, we first compare our deep instruction tuning (DIT) methods with the default fusion scheme of SAM and its alternatives, i.e., adding more layers to the mask decoder. We can see that the default fusion scheme performs much worse than DITs on RefCOCO under the setting of either fully tuning or freezing backbone, suggesting that the default lightweight decoder is insufficient in text following. With more decoding layers, this issue can be alleviated to some extent, but the performance still lags behind DITs especially when the image encoder is frozen. To explain, the additional fusion layers can enhance the cross-modal interactions, but also undermines the powerful segmentation capability of SAM pre-trained on massive data. In contrast, without modifying the main structure of SAM, our DIT methods can improve the performance greatly. Another interesting finding from Tab. LABEL:Ablation_of_Tuning_method is that the advantage of layer-wise DIT is more obvious under the setting of freezing backbone. This result confirms our assumption that layer-wise DIT can better help to project text tokens onto the semantic space of SAM. Overall, Tab. LABEL:Ablation_of_Tuning_method well validates the effectiveness of DITs for text-instructed SAM.
4.3.2. Combination with prompt tuning
@@
Tab. LABEL:framwork present the results of combing DITs with conventional the prompt tuning (Zhou et al., 2022a, b; Wu et al., 2023b), i.e., adding learnable tokens to the text prompts. Under end-to-end tuning, the addition of learnable tokens can improve performance to a certain extent, mainly due to the increase of the parameter capacity of DIT for better text tuning. However, these prompt tokens are not that helpful for layer-wise DIT, on the contrary, reducing performance greatly. This result may suggests that the text words are well projected onto each layer of SAM with independent projectors. The combination with parameterized tokens somewhat declines their semantics. These results also indicate the difference between DITs and prompt tuning, i.e., one is for semantic projection while the other is for task adaption.
4.3.3. The impact of different text encoder on Layer-wise DIT
Tab. LABEL:text_encoder examines the impact of different language encoders. It can be seen that BERT outperforms CLIP in most cases, though the gap between the two encoders is not significant. For example, CLIP performs slightly better than BERT in terms of oIoU on testB. We believe this is due to the typically concise descriptions in the dataset, resulting in a relatively limited impact of the text encoder on DIT-SAM.
4.3.4. The effect of freezing encoders